# Upper bounds of focusing light through multimode fibers

### v1

@@ III The enhancement factor and its practical limitations
We present a general form of the enhancement factor at the output $m$ in terms of two major parameters; the input participation ratio $R$ and the phase error coefficient $\Phi$:
@@
where $\alpha\equiv\left\langle A_{n}\right\rangle_{n}^{2}/\left\langle A_{n}^{2}\right\rangle_{n}$ defines the spatial homogeneity of the amplitude profile incident on the SLM, $R_{m}\equiv\left\langle|t_{mn}|\right\rangle_{n}^{2}/\left\langle|t_{mn}|^{2}\right\rangle_{n}$ quantifies the input participation ratio at the output $m$, $\Phi_{m}\equiv\left\langle{\rm cos}(\delta\phi_{mn})\right\rangle_{n}^{2}$ quantifies the phase error coefficient at the output $m$, and $N$ represents the number of controlled degrees of freedom at the input wavefront. Here $\left\langle\right\rangle_{n}$ denotes averaging over inputs $n$. The phase error coefficient $\Phi$ varies between 0 and 1, where, in the absence of phase errors, it equals $\Phi=1$ van Putten (2011). In our experiments, $\alpha=1$, since we illuminate the SLM with a flat top laser beam expanded from a single-mode fiber output. In our experimental system, the input index $n$ corresponds to the SLM segment index (approximately the Fourier components at the proximal end) and the output index $m$ to the CCD camera pixels (position at the distal end), respectively. Note that the enhancement factor is defined differently in certain studies Gomes et al. (2022); Duan et al. (2023), whereas the enhancement factor we establish here is standardized Vellekoop and Mosk (2007, 2008b, 2008a); Papadopoulos et al. (2012); Yılmaz et al. (2013); Vellekoop (2015); Amitonova et al. (2016); Resisi et al. (2020); Mastiani et al. (2024) and possesses a clearly defined upper limit $\eta=N_{\rm mode}$ with complete ($N=N_{\rm mode}$) full-field (amplitude and phase) input modulation.
III.1 The enhancement factor for phase-only input modulation
@@
In phase-only wavefront shaping experiments, evaluating the input participation ratio $R_{m}$ is essential, as it quantifies the fraction of input degrees of freedom that effectively contribute to the output at position $m$ under phase-only modulation. We quantify $R_{m}$ both in experiments and numerical simulations. To obtain the numerical participation ratio, we first compute the transmission matrix using the mode decomposition method Lee (2023) (see Section II in the supplementary material for details). To evaluate the participation ratio in the desired input basis, we apply a basis transformation to the columns of the transmission matrix and compute the participation ratio in the transformed basis. To obtain the experimental input participation ratio, we complemented phase measurements of the transmission matrix—acquired using the dual reference algorithm in the Hadamard basis—with amplitude measurements by coupling only the signal component of the light from the SLM into the fiber. This approach allowed us to reconstruct both the amplitude and phase of the transmission matrix elements $t_{mn^{\prime}}$, where $n^{\prime}$ represents the Hadamard vector index. We then applied a Hadamard transform as $t_{mn}=\sum_{n^{\prime}=1}^{N}t_{mn^{\prime}}H_{n^{\prime}n}$ to convert the transmission matrix into the canonical (SLM pixel) basis $n$. Here $H_{n^{\prime}n}$ represents the unitary Hadamard transform matrix. The participation ratio was subsequently calculated on this basis, as our phase-only modulation experiments are performed in canonical bases.
@@
Fig. 2(a) presents both numerical and experimental participation ratios $R$ as a function of the normalized radial distance $r/a$. The numerical results in the Fourier basis closely align with the experimental data, exhibiting the same overall trend. Notably, when phase-only input modulation is performed on the Fourier basis, $R$ remains constant regardless of the focal position at the distal end of the fiber. Additionally, placing the spatial light modulator (SLM) in the Fourier plane of the MMF’s proximal end yields a participation ratio of $R=\pi/4$, consistent with the well-known value observed in phase-only wavefront shaping through random media van Beijnum (2009). In the experiments, the number of guided modes is $N_{\rm mode}=15,178$ for a fiber with a core radius of $a=100$ µm and numerical aperture (NA) of 0.22. In the simulations, we consider four fibers with core radii of $a$ = 15 µm, 22 µm, 28 µm, and 38 µm (all with NA = 0.22), corresponding to $N_{\rm mode}$ = 180, 374, 606, and 1,114, respectively. All numerical values are calculated for light with a wavelength of 561 nm.
@@
While implementing phase-only modulation directly on the fiber mode basis is not practical, it offers valuable insights into the fundamental behavior of phase-only wavefront shaping. For this reason, we also numerically calculated the participation ratio on the fiber mode basis. Our simulations reveal that when phase-only modulation is applied on this basis, the participation ratio $R$ shows a strong dependence on the radial position of the focal point at the distal end of the MMF. At the center of the fiber core, the participation ratio drops below 0.2, as most fiber mode wavefunctions exhibit ring-shaped profiles and contribute minimally at the core center.
@@
Fig. 2(b) shows the radially averaged theoretical and experimental participation ratios $R$ (both for Fourier basis) as a function of the normalized number of degrees of freedom $N/N_{\text{mode}}$, where the experimental $N_{\text{mode}}$ is determined from the equation $V=2\pi a\text{NA}/\lambda$, with $\lambda=561$ nm, $a=100$ µm (core radius), and NA = 0.22. For our fiber, the number of modes per polarization is $N_{\text{mode}}=V^{2}/4=15,178$. Our experimental and numerical results consistently show that when phase-only input modulation is performed on the Fourier basis, the participation ratio remains $R=\pi/4$, regardless of the number of input degrees of freedom.
@@
Fig. 2(c) presents the enhancement factor $\eta$, averaged over radial ($r$) and azimuthal ($\theta$) coordinates, as a function of $N$. We observe that the experimental enhancement factor closely approaches the theoretical limit without phase errors, $\eta=(\pi/4)(N-1)+1$, when the transmission matrix is measured on the Hadamard basis. For $N=2,000$, the enhancement factor obtained in the canonical basis is lower than the theoretical prediction, as the signal-to-noise ratio decreases with increasing $N$, leading to phase errors ($\Phi<1$).
@@
Our theoretical and experimental results demonstrate that the widely recognized $\pi/4$ factor observed in wavefront shaping through disordered media also holds for wavefront shaping through multimode fibers. This similarity arises because each degree of freedom in the Fourier space couples randomly to multiple superpositions of fiber modes, leading to output speckle statistics that closely resemble those seen in disordered media.
III.2 The effect of phase errors on the enhancement factor
@@
Accurate transmission matrix measurements are essential to approach the theoretical upper bound of the enhancement factor. However, experimental noise introduces phase errors that degrade these measurements Vellekoop and Mosk (2008a); Yılmaz et al. (2019). Here we introduce a predictive methodology to quantify these phase errors by applying the phase-shift interferometry measurement twice for the same SLM input pattern. In the absence of phase errors, the measured phases $\phi_{mn}^{(1)}$ and $\phi_{mn}^{(2)}$ must be identical for the same input $n$ and output $m$. However, in practice, experimental noise gives rise to phase errors, $\delta\phi_{mn}^{(1)}$ and $\delta\phi_{mn}^{(2)}$, in both measurements. We extract these phase errors as $\delta\phi_{mn}^{\prime}=\arg\left(e^{i(\phi_{mn}^{(2)}-\phi_{mn}^{(1)})}\right)$. A sketch is shown in Fig. 3(a) to describe the phase error measurement concept. Assuming the phase errors follow a Gaussian distribution with zero mean—which is a reasonable assumption as evidenced by Figs. 3(c) and Figs. 3(d)—we estimate the phase errors $\delta\phi_{mn}$ in a single four phase shift interferometry measurement (see section I-D in the supplementary material for details).
@@
Fig. 3(b) presents the standard deviation $\sigma_{\delta\phi^{\prime}}$ of the least-squares fitted Gaussian functions to the experimental phase error histograms for different numbers of degrees of freedom $N$. The Hadamard basis maintains a consistently low $\sigma_{\delta\phi^{\prime}}$, indicating that phase errors remain minimal variations. In contrast, the canonical basis exhibits an increasing $\sigma_{\delta\phi^{\prime}}$ as $N$ grows, revealing a broadening phase error distribution. This trend suggests that phase errors in the canonical basis become more pronounced with an increasing number of degrees of freedom, leading to greater deviations from the ideal constructive interference.
@@
Wavefront shaping relies on precise constructive interference at the target position $m$, which makes it sensitive to experimental noise that introduces phase errors. To characterize the level of constructive interference in our experiments, we introduce the phase error coefficient $\Phi_{m}\equiv\left\langle{\rm cos}(\delta\phi_{mn})\right\rangle_{n}^{2}$ at the target position $m$. This coefficient equals unity in the absence of phase errors and approaches zero when the phase measurements contain no usable information. Fig. 4(a) illustrates how phase deviations $\delta\phi_{mn}$ disrupt the alignment of field contributions, reducing $\Phi$ at the output position $m$. In an ideal wavefront shaping experiment, the phases of the individual field contributions are perfectly aligned, leading to strong constructive interference and high intensity at the target. In practical, high-quality experiments, small phase errors reduce this alignment slightly, resulting in $\Phi$ values close to 1. However, in canonical basis measurements, $\Phi$ is generally much lower—especially for large $N$—due to more significant phase mismatches.
@@
To quantify the impact of accumulated phase errors, we experimentally quantified the phase error coefficient $\Phi$ at each output position $m$ and computed its ensemble-averaged value $\langle\Phi_{mn}\rangle_{n}$ as a function of the number of degrees of freedom $N$, as shown in Fig. 4(b). In the Hadamard basis, $\Phi$ remains close to 1, while in the canonical basis, $\Phi$ decreases significantly as $N$ increases. This degradation is due to the fact that Hadamard wavefront shaping maintains a high signal-to-noise ratio by balancing the signal and reference fields, ensuring maximum interferometric visibility. In contrast, on the canonical basis, as $N$ increases, the signal diminishes, reducing the signal-to-noise ratio, degrading interference visibility, and limiting the accuracy of the measured transmission matrix.
@@
To further explore how the photon budget influences phase errors, we examine the behavior of $\Phi$ for a fixed $N=8000$ as a function of the incident power on the CCD camera (seen in Fig. 4(c)). The incident power is adjusted by placing neutral density (ND) filters in front of the CCD chip. The Hadamard basis maintains $\Phi\approx 1$ even as the available photon budget decreases, demonstrating its robustness to photon-limited conditions Jang et al. (2017). This stability arises because Hadamard-based measurements preserve high interferometric contrast and balanced signal-to-reference ratios, maintaining a reliable signal-to-noise ratio for phase retrieval. However, on the canonical basis, $\Phi$ steadily drops as the photon budget decreases. The degradation stems from a reduced signal-to-noise ratio, leading to increased phase errors. As expected, when the signal photon count approaches zero, phase error coefficient estimation deteriorates for both bases, ultimately preventing phase retrieval.
@@ IV Results and Discussion
Figure 5: The mean enhancement factor $\eta$, averaged over the azimuthal $\theta$ positions and radial distance $r$ at the fiber’s distal end, as a function of the number of degrees of freedom $N$. Solid blue and red lines show the theoretical enhancement factors, while dashed lines represent experimental measurements for wavefront shaping on the canonical and Hadamard bases, respectively. Error bars indicate the standard deviation of enhancement factors measured at different focal positions on the fiber’s distal end.
@@
Here, we report the enhancement factor values for focusing at various radial distances ($r$) and azimuthal angles ($\theta$) within the fiber core at the distal end, for different values of $N$, using transmission matrices measured in both the canonical and Hadamard bases. Additionally, we provide the predicted enhancement factors obtained from our method, which uses the measured phase error coefficients $\Phi$ for both bases.
@@
In Fig. 5, we display the measured and predicted enhancement factor $\eta$ as a function of the number of degrees of freedom, $N$. Transmission matrix measurements in the Hadamard basis result in improved enhancement factors, thanks to its superior signal-to-noise ratio, particularly as $N$ increases. However, for $N<952$, the results from both the canonical and Hadamard bases are nearly identical, indicating that the advantages of the Hadamard basis become more pronounced as the number of degrees of freedom grows.
@@
For lower values of $N$ ($N<3000$), the measured enhancement factor closely aligns with the expected theoretical values. For larger $N$, however, the measured enhancement deviates from the theoretical predictions, consistently remaining lower due to experimental limitations. Below $N=3{,}000$, phase errors caused by experimental noise are the primary limiting factor. Beyond this threshold ($N>3{,}000$), transmission matrix decorrelation becomes the dominant constraint, especially during measurement periods extending up to four hours.
V Conclusion

### v3

@@ III The enhancement factor and its practical limitations
We present a general form of the enhancement factor at the output $m$ in terms of two major parameters: the input participation ratio $R$ and the phase-error coefficient $\Phi$:
@@
where $\alpha\equiv\left\langle A_{n}\right\rangle_{n}^{2}/\left\langle A_{n}^{2}\right\rangle_{n}$ defines the spatial homogeneity of the amplitude profile incident on the SLM, $R_{m}\equiv\left\langle|t_{mn}|\right\rangle_{n}^{2}/\left\langle|t_{mn}|^{2}\right\rangle_{n}$ quantifies the input participation ratio at the output $m$, $\Phi_{m}\equiv\left\langle{\rm cos}(\delta\phi_{mn})\right\rangle_{n}^{2}$ quantifies the phase-error coefficient at the output $m$, and $N$ represents the number of controlled degrees of freedom at the input wavefront. Here $\left\langle\right\rangle_{n}$ denotes averaging over inputs $n$. The phase error coefficient $\Phi$ varies between 0 and 1, where, in the absence of phase errors, it equals $\Phi=1$ [55]. In our experiments, $\alpha=1$, since we illuminate the SLM with a flat-top laser beam expanded from a single-mode fiber output. In our setup, each input degree of freedom, indexed by $n$, corresponds to a $6\times 6$ pixel segment on the SLM (approximately representing a single Fourier component at the proximal end), while each output index $m$ corresponds to a CCD camera pixel, imaging the distal end.
The enhancement factor used in this work is defined as the ratio between the intensity at the target position $m$ after wavefront shaping and the average intensity at the same position $m$ measured over many independent random input wavefronts. In some studies, however, the enhancement factor is instead defined as the ratio between the focused intensity and the average background intensity after focusing—an approach that effectively serves as a proxy for the signal-to-noise ratio (SNR) [40, 56, 27].
When the number of controlled degrees of freedom $N$ is much smaller than the total number of modes $M$ ($N\ll M$), these two definitions are equivalent. As $N$ increases, however, non-Gaussian contributions to the intensity correlations become significant. Long-range correlations, for instance, can raise the background intensity level, as has been experimentally demonstrated for wavefront shaping and focusing through disordered media [57, 58]. This makes the two definitions diverge in the non-Gaussian regime.
Multimode fibers (MMFs) represent a distinct case—particularly when nearly all propagating modes can be controlled ($N\approx M$) and the transmission matrix is nearly unitary. In the ideal limit of full-field (amplitude and phase) control, focusing is exact, and the SNR diverges as $N\to M$, being limited only by practical constraints rather than theoretical bounds.
In this work, we adopt the standardized definition of the enhancement factor [31, 57, 44, 3, 45, 19, 35, 47, 43], which has a well-defined upper bound of $\eta=M$ for complete ($N=M$) full-field input modulation and $\eta=R(N-1)+1$ for phase-only input modulation (see Appendix B for details).
III.1 Enhancement factor for phase-only input modulation
@@
In phase-only wavefront shaping experiments, evaluating the input participation ratio $R_{m}$ is essential, as it quantifies the fraction of input degrees of freedom that effectively contribute to the output at position $m$ under phase-only modulation. We quantify $R_{m}$ both in experiments and numerical simulations. To obtain the numerical participation ratio, we first compute the transmission matrix using the mode decomposition method [59] (see Appendix B for details). To evaluate the participation ratio in the desired input basis, we apply a basis transformation to the columns of the transmission matrix and compute the participation ratio in the transformed basis. To obtain the experimental input participation ratio, we complemented phase measurements of the transmission matrix—acquired using the dual reference algorithm in the Hadamard basis—with amplitude measurements by coupling only the signal component of the light from the SLM into the fiber. This approach allowed us to reconstruct both the amplitude and phase of the transmission-matrix elements $t_{mn^{\prime}}$, where $n^{\prime}$ represents the Hadamard vector index. We then applied a Hadamard transform as $t_{mn}=\sum_{n^{\prime}=1}^{N}t_{mn^{\prime}}H_{n^{\prime}n}$ to convert the transmission matrix into the canonical (SLM pixel) basis $n$. Here $H_{n^{\prime}n}$ represents the unitary Hadamard transform matrix. The participation ratio was subsequently calculated on this basis, as our phase-only modulation experiments are performed in canonical bases.
@@
Fig. 2(a) presents both numerical and experimental participation ratios $R$ as a function of the normalized radial distance $r/a$ within the core at the distal end. The numerical results in the Fourier basis closely align with the experimental data, exhibiting the same overall trend. Notably, when phase-only input modulation is performed on the Fourier basis, $R$ remains constant regardless of the focal position at the distal end of the fiber. Additionally, placing the spatial light modulator (SLM) on the Fourier plane of the MMF’s proximal end yields a participation ratio of $R=\pi/4$, consistent with the well-known value observed in phase-only wavefront shaping through disordered media [51]. In the experiments, the number of guided modes is $M=15,178$ for a fiber with a core radius of $a=100$ µm and numerical aperture (NA) of 0.22. In the simulations, we consider four fibers with core radii of $a$ = 15 µm, 22 µm, 28 µm, and 38 µm (all with NA = 0.22), corresponding to $M$ = 180, 374, 606, and 1,114, respectively. All numerical values are calculated for light with a wavelength of 561 nm.
@@
While implementing phase-only modulation directly on the fiber mode basis is not practical, it offers valuable insights into the fundamental behavior of phase-only wavefront shaping. For this reason, we also calculated the participation ratio numerically on a fiber mode basis. Our simulations reveal that when phase-only modulation is applied on this basis, the participation ratio $R$ shows a strong dependence on the radial distance of the focal point at the distal end of the MMF. At the center of the fiber core, the participation ratio drops below 0.2, as most fiber mode wavefunctions exhibit ring-shaped profiles and contribute minimally at the core center.
@@
Fig. 2(b) shows the radially averaged theoretical and experimental participation ratios $R$ (both for Fourier basis) as a function of the normalized number of degrees of freedom $N/M$, where the experimental $M$ is determined from the equation $V=k_{0}a\text{NA}$, with $k_{0}=2\pi/\lambda$, $\lambda=561$ nm, $a=100$ µm (core radius), and NA = 0.22. For our fiber, the number of modes per polarization is $M=V^{2}/4=15,178$. Our experimental and numerical results consistently show that when phase-only input modulation is performed on the Fourier basis, the participation ratio remains $R=\pi/4$, regardless of the number of input degrees of freedom.
@@
Fig. 2(c) presents the enhancement factor $\eta$, averaged over radial ($r$) and azimuthal ($\theta$) positions, as a function of $N$. We observe that the experimental enhancement factor closely approaches the theoretical limit without phase errors, $\eta=(\pi/4)(N-1)+1$, when the transmission matrix is measured on the Hadamard basis. For $N=2,000$, the enhancement factor obtained in the canonical basis is lower than the theoretical prediction, as the signal-to-noise ratio decreases with increasing $N$, leading to phase errors ($\Phi<1$).
@@
Our theoretical and experimental results show that the widely recognized $\pi/4$ factor in wavefront shaping through disordered media also appears in step-index multimode fibers when phase-only modulation is applied at the Fourier plane of the fiber’s proximal end. Numerical simulations further indicate that this $\pi/4$ factor persists for graded-index multimode fibers under the same modulation scheme, except near the core edge (see Appendix B). This correspondence arises because each Fourier-space degree of freedom couples randomly to multiple superpositions of fiber modes, yielding output speckle statistics that closely resemble those of disordered media.
Our framework for predicting the enhancement factor $\eta$ via the participation ratio $R$ is general and applies to any linear optical medium for which a transmission matrix can be obtained. Whether the medium is a scattering slab, a chaotic cavity, a step-index MMF, or a graded-index MMF, the theory holds as long as the transmission matrix accurately describes the input–output field relationship (see Appendix B for details).
III.2 Effect of phase errors on enhancement factor
@@
Accurate transmission-matrix measurements are crucial for approaching the theoretical upper bound of the enhancement factor. However, experimental noise introduces phase errors that degrade these measurements [44, 53]. Here, we introduce a predictive methodology to quantify these phase errors by applying the phase-shift interferometry measurement twice for the same SLM input pattern. In the absence of phase errors, the measured phases $\phi_{mn}^{(1)}$ and $\phi_{mn}^{(2)}$ must be identical for the same input $n$ and output $m$. However, in practice, experimental noise gives rise to phase errors, $\delta\phi_{mn}^{(1)}$ and $\delta\phi_{mn}^{(2)}$, in both measurements. We extract these phase errors as $\delta\phi_{mn}^{\prime}=\arg\left(e^{i(\phi_{mn}^{(2)}-\phi_{mn}^{(1)})}\right)$. A sketch is shown in Fig. 3(a) to describe the phase-error measurement concept. Assuming the phase errors follow a Gaussian distribution with zero mean—which is a reasonable assumption as evidenced by Figs. 3(c) and Figs. 3(d)—we estimate the phase errors $\delta\phi_{mn}$ in a single four-phase shift interferometry measurement (see Appendix A for details).
@@
Fig. 3(b) presents the standard deviation $\sigma_{\delta\phi^{\prime}}$ of the least-squares fitted Gaussian functions to the experimental phase-error histograms for different numbers of degrees of freedom $N$. The Hadamard basis maintains a consistently low $\sigma_{\delta\phi^{\prime}}$, indicating that phase errors remain minimal variations. In contrast, the canonical basis exhibits an increasing $\sigma_{\delta\phi^{\prime}}$ as $N$ increases, revealing a broadening phase-error distribution. This trend suggests that phase errors in the canonical basis become more pronounced as the number of degrees of freedom increases, leading to greater deviations from ideal constructive interference.
@@
Wavefront shaping relies on precise constructive interference at the target position $m$, which makes it sensitive to experimental noise that introduces phase errors. To characterize the level of constructive interference in our experiments, we introduce the phase-error coefficient $\Phi_{m}\equiv\left\langle{\rm cos}(\delta\phi_{mn})\right\rangle_{n}^{2}$ at the target position $m$. This coefficient equals unity in the absence of phase errors and approaches zero when the phase measurements contain no usable information. Fig. 4(a) illustrates how phase deviations $\delta\phi_{mn}$ disrupt the alignment of field contributions, reducing $\Phi$ at the output position $m$. In an ideal wavefront shaping experiment, the phases of the individual field contributions are perfectly aligned, leading to strong constructive interference and high intensity at the target. In practical, high-quality experiments, small phase errors reduce this alignment slightly, resulting in $\Phi$ values close to 1. However, in canonical basis measurements, $\Phi$ is generally much lower—especially for large NN—due to more significant phase mismatches.
@@
To quantify the impact of accumulated phase errors, we experimentally quantified the phase-error coefficient $\Phi$ at each output position $m$ and computed its ensemble-averaged value $\langle\Phi_{mn}\rangle_{n}$ as a function of the number of degrees of freedom $N$, as shown in Fig. 4(b). In the Hadamard basis, $\Phi$ remains close to 1, while in the canonical basis, $\Phi$ decreases significantly as $N$ increases. Hadamard-based dual-reference transmission-matrix measurement maintains a high signal-to-noise ratio by balancing the signal and reference fields, ensuring maximum interferometric visibility. In contrast, on the canonical basis, as $N$ increases, the signal diminishes, reducing the signal-to-noise ratio, degrading the visibility of interference, and limiting the accuracy of the measured transmission matrix.
Phase singularities—points of zero intensity in the reference speckle patterns that lead to undefined phases—can, in principle, compromise the accuracy of common-path interferometry–based transmission-matrix measurements. In our setup, speckles are sampled well above the Nyquist criterion (more than three CCD pixels per speckle grain), ensuring that all pixels receive non-zero intensity and effectively mitigating this issue. Moreover, our Hadamard-based dual-reference measurement scheme balances the signal and reference, further enhancing the robustness of phase retrieval. As a result, null phases do not degrade the measured enhancement factor. For comparison, in sequential stepwise algorithms, phase singularities can introduce phase errors; however, such effects are inherently captured by the phase-error coefficient $\Phi$, which quantifies all sources of phase inaccuracy.
@@
To further explore how the photon budget influences phase errors, we examine the behavior of $\Phi$ for a fixed $N=8000$ as a function of the incident power on the CCD camera (seen in Fig. 4(c)). The incident power is adjusted by placing neutral density (ND) filters in front of the CCD chip. The Hadamard basis maintains $\Phi\approx 1$ even as the available photon budget decreases, demonstrating its robustness to photon-limited conditions [60]. This stability arises because Hadamard-based measurements preserve high interferometric contrast and balanced signal-to-reference ratios, maintaining a reliable signal-to-noise ratio for phase retrieval. However, on the canonical basis, $\Phi$ steadily drops as the photon budget decreases. The degradation stems from a reduced signal-to-noise ratio, leading to increased phase errors. As expected, when the signal photon count approaches zero, phase-error coefficient estimation deteriorates for both bases, ultimately preventing phase retrieval.
@@ IV Results and Discussion
Figure 5: The mean enhancement factor $\eta$, averaged over the azimuthal $\theta$ position and radial distance $r$ at the fiber’s distal end, as a function of the number of degrees of freedom $N$ is shown. Solid blue and red lines show the theoretical enhancement factors, while dashed lines represent experimental measurements for wavefront shaping on the canonical and Hadamard bases, respectively. Error bars indicate the standard deviation of enhancement factors measured at different focal positions on the fiber’s distal end.
@@
Here, we report the enhancement factor values for focusing at various radial distances ($r$) and azimuthal angles ($\theta$) within the fiber core at the distal end, for different values of $N$, using transmission matrices measured in both the canonical and Hadamard bases. Additionally, we provide the predicted enhancement factors obtained from our method, which uses the measured phase-error coefficients $\Phi$ for both bases.
@@
In Fig. 5, we display the measured and predicted enhancement factor $\eta$ as a function of the number of degrees of freedom, $N$. Transmission-matrix measurements in the Hadamard basis result in improved enhancement factors, thanks to its superior signal-to-noise ratio, particularly as $N$ increases. However, for $N<952$, the results from both the canonical and Hadamard bases are nearly identical, indicating that the advantages of the dual-reference method in the Hadamard basis become more pronounced as the number of degrees of freedom grows.
@@
For lower values of NN ($N<3000$), the measured enhancement factor closely aligns with the expected theoretical values. For larger $N$, however, the measured enhancement deviates from the theoretical predictions, consistently remaining lower due to experimental limitations. Below N=3, [^0] N=3{,}000, phase errors caused by experimental noise are the primary limiting factor. Beyond this threshold (N>3, [^0] N>3{,}000), transmission-matrix decorrelation becomes the dominant constraint, especially during measurement periods extending up to four hours.
V Conclusions